OpenStudy (jazzyluvsu123):

(-3)-x=12 MEDAL & Fan

OpenStudy (skullpatrol):

Any ideas? :-)

OpenStudy (anonymous):

answer is -15

OpenStudy (jazzyluvsu123):

15 or - 15

OpenStudy (skullpatrol):

Do you know how to work with negative numbers in solving equations?

OpenStudy (jazzyluvsu123):

sorta im not the best at it

OpenStudy (skullpatrol):

Add 3 to both sides, what do you get?

OpenStudy (jazzyluvsu123):

-x + 3 = 12?

OpenStudy (skullpatrol):

(-3) - x = 12 3 + (-3) - x = 3 + 12

OpenStudy (skullpatrol):

Can you simply that^